# Transitions to Intermittent Chaos in Quorum Sensing Dynamics

### v2

@@ 6 Concluding Remarks
This present study has investigated the profound influence of distinct time delays on the dynamical behaviour of a QS-inspired system, employing a nonlinear delay differential equation framework. Specifically, we modelled a system comprising motile and static bacterial subpopulations, each exhibiting distinct response times to autoinducer signals, thereby introducing two independent delay parameters, $\tau_{1}$ and $\tau_{2}$. This approach acknowledges the inherent temporal lags associated with the intricate network of biochemical processes underpinning QS, which may include signal production, transport diffusion features, and cellular response.
@@
Our analysis commenced explicitly incorporating two distinct delays, building upon a well-established activator-inhibitor framework to effectively represent the dynamic interplay between motile (activator) and static (inhibitor) bacterial populations. Subsequently, we conducted a detailed analysis of the existence and local stability properties of the system’s steady-states. Upon deriving the pseudo-characteristic polynomial and meticulously examining the conditions governing the emergence of purely imaginary eigenvalues, we established the potential for Hopf bifurcations as the delay parameters were slowly varied. Our findings show that the presence of multiple delays, particularly when characterised by significant disparities in magnitude, can dramatically alter the system’s stability features and promote the emergence of complex nonlinear oscillatory behaviour. Specifically, we have demonstrated that the steady-states $E_{1}$ and $E_{2}$ can undergo Hopf bifurcations, losing stability as the difference between the delays, $h=\tau_{1}-\tau_{2}$, is increased, while the steady-state $E_{3}$ robustly maintains its stability across the explored range of delay values. The observed difference in the magnitude of $h$ required to induce instability in $E_{1}$ compared to $E_{2}$ suggests a complex and nuanced interplay between the two distinct delays and the inherent dynamical properties of each respective steady-state.
The results presented in this study make a significant contribution to our fundamental understanding of the role of time delays in shaping the dynamics of QS systems. Upon explicitly incorporating distinct delays for different state-components, we have shown how temporal factors can dramatically influence system stability and give rise to a spectrum of complex dynamical behaviours, including intermittent chaos. These findings offer crucial contributions to our understanding of microbial communication and have the potential to inform the development of innovative strategies for targeted manipulation of QS systems. Future research directions could explore an extension of the current model to incorporate spatial heterogeneity and more complex interaction topologies, important factors that are beyond the scope of the present analysis and will be addressed in future work. Moreover, the insights gained from this study may prove invaluable in the design and engineering of synthetic biological systems where precise control of gene expression and population dynamics is paramount.
In summary, we have exhibited the impact of heterogeneous time delays on QS dynamics. Upon modelling subpopulation-specific response times, we have revealed the complex interplay between temporal heterogeneity, stability, and emergent oscillations. These findings promote our understanding of delay-mediated regulation and may offer a foundation for developing control strategies in synthetic biology, emphasising the necessity of considering temporal heterogeneity in certain complex biological systems.
@@
Moreover, the presented investigation has yielded key insights into the temporal dynamics of QS; however, the inherent spatial heterogeneity arising from transport phenomena constitutes a critical determinant of system behaviour. To achieve a more comprehensive and predictive understanding, future research will extend the current model to explicitly incorporate transport dynamics. This integration is essential for capturing the spatiotemporal complexity inherent in QS systems.
@@ Acknowledgements
AF thanks the financial support by ‘Programa CAPSEM I+DT’, Faculty of Engineering, UNAM. MAGO would like to thank Colegio de Ciencia y Tecnología for its financial support through project UACM CCYT2023-IMP-05 and CONAHCyT-SNII. VFBM thanks the financial support by Asociación Mexicana de Cultura A.C.

### v3

@@ 6 Concluding Remarks
This study delves into the influence of varied time delays on the dynamic behaviour of a QS-inspired system, employing a nonlinear DDE framework. Specifically, we model a system with motile and static bacterial sub-populations, each potentially exhibiting distinct response times to autoinducer signals. This necessitates the incorporation of two independent delay parameters, $\tau_{1}>0$ and $\tau_{2}>0$, which represent the intrinsic temporal lags associated with the intricate biochemical processes underlying QS, such as signal production and cellular response.
@@
Our analysis begins by explicitly introducing these two distinct delays into a well-established activator-inhibitor-like framework, effectively representing the dynamic interplay between motile (activator) and static (inhibitor) bacterial sub-populations. We then conduct a thorough analysis of the existence and local stability properties of the system’s steady-states. Upon deriving the pseudo-characteristic polynomial and examining the conditions for purely imaginary eigenvalues, we identify the potential for Hopf bifurcations as the delay parameters are slowly varied. Our findings indicate that the presence of multiple delays, particularly when their magnitudes differ significantly, can substantially alter the system’s stability and promote the emergence of complex nonlinear oscillatory behaviour. We observe that steady-states $E_{1}$ and $E_{2}$ can undergo Hopf bifurcations, potentially losing stability as the difference between the delays, $h=\tau_{1}-\tau_{2}$, increases. In contrast, steady-state $E_{3}$ tends to maintain its stability across the explored range of delay values. The observed disparity in the magnitude of $h$ required to induce instability in $E_{1}$ versus $E_{2}$ suggests a complex and intricate interaction between the two distinct delays and the inherent dynamic properties of each steady-state.
Our results underscore the critical importance of considering the inherent temporal dynamics within biological QS systems for accurate modelling. Upon incorporating distinct time delays for motile and static bacteria—representing the characteristic timescales of signal production and cellular response—we reveal their influence on system behaviour. The observed instabilities, the potential for oscillatory behaviour via Hopf bifurcations, and the varied stability of steady-states $E_{1}$, $E_{2}$, and $E_{3}$ under different delay conditions emphasise that the intrinsic timescales of biological processes are crucial factors governing QS dynamics. The difference in delays, characterised by $h$, emerges as a key parameter influencing these dynamics, highlighting the necessity of accounting for temporal heterogeneity in models of microbial communication to capture realistic behaviours.
The analysis presented here offers novel insights into the role of time delays in shaping the dynamics of QS systems. Upon explicitly incorporating distinct delays for different state components, we illustrate how temporal factors can significantly influence system stability and potentially give rise to a range of complex dynamic behaviours, including intermittent chaos. The emergence of intermittent chaos in our QS-inspired model, particularly when compared to existing literature on chaotic dynamical features in similar systems (briefly addressed in Section 1), represents a fairly novel route to chaos in such systems.
Intermittent chaos has been a rare and recent discovery within the realm of QS models. While a limited number of studies have explicitly identified and explored this particular dynamical regime, the work by Zakharova et al. [59] demonstrated its presence in a model of bacterial QS, linking its emergence to specific parameter ranges governing signal production and reception. Complementing this, Singh et al. [49] similarly reported intermittent chaotic dynamics within a different QS model, highlighting its potential implications for understanding the robustness and adaptability of bacterial communication networks.
A key distinction between these two foundational works lies in their focus. In [49], the authors focused on the general emergence of intermittent chaotic behaviour from the inherent nonlinearities and feedback mechanisms within a broad QS framework, showcasing the possibility of such complex dynamics without pinpointing a specific driving parameter. In contrast, the results reported in [59] specifically gathers the role of time delays, demonstrating how these temporal lags, characteristic of biochemical processes in QS, may act as a crucial bifurcation parameter directly leading to the onset of intermittent chaos. Thus, while both contribute significantly to the understanding of chaotic features in QS, the latter offers a more direct mechanistic link between a specific biological timescale and the appearance of intermittent chaos.
In the present work, we further explore this phenomenon by revealing its emergence in QS-inspired systems with heterogeneous temporal delays. This represents a more complex scenario of temporal heterogeneity, demonstrating how the interplay of multiple, distinct delays can lead to intermittent chaos in QS-inspired systems. Our analysis therefore expands the understanding of intermittent chaos in QS dynamics, moving beyond single-delay mechanisms to consider more biologically realistic scenarios with varied response times.
These studies collectively underscore that intermittent chaos remains a less explored area within QS dynamics, representing a significant avenue for future research. This suggests that intermittent chaos may play a crucial role in the flexibility and adaptability of bacterial communication, particularly in fluctuating environments, and warrants further in-depth investigation.
Our findings suggest that identifying specific “chaotic windows” within the parameter space of varying delay values can be quite challenging. Even small changes in either $\tau_{1}$ or $\tau_{2}$ can lead to relatively abrupt transitions between stable, oscillatory, and chaotic regimes, highlighting the system’s sensitivity to its initial conditions and the intricate interplay between the delays. These findings may provide useful insights into microbial communication and could potentially inform the development of innovative strategies for targeted manipulation of QS systems. Moreover, the insights gained from this study may prove helpful in the design and engineering of synthetic biological systems where precise control of gene expression and population dynamics is crucial.
@@
While this investigation has provided key insights into the temporal dynamics of QS, the inherent spatial heterogeneity arising from transport phenomena represents an essential aspect of system behaviour. Future research could explore extending the current model to incorporate such spatial heterogeneity within complex interaction topologies. Although these are important factors beyond the scope of this current analysis, their integration is essential for achieving a more comprehensive understanding, thus capturing the spatio-temporal complexity observed in QS systems.
@@ Acknowledgements
The authors appreciate the reviewers’ valuable comments and insightful suggestions. AFP thanks the financial support by ‘Programa CAPSEM I+DT’, Faculty of Engineering, UNAM. MAGO would like to thank Colegio de Ciencia y Tecnología for its financial support through project UACM CCYT2023-IMP-05 and CONAHCyT-SNII. VFBM thanks the financial support by Asociación Mexicana de Cultura A.C.